5kg of sugar and 7kg of rice costs ₹153 & 7kg of sugar & 5kg of rice costs ₹147 how much would 6kg of sugar & 10kg of rice costs simultaneously step by step
Answer by Cromlix(4381)

Hi there,
Make sugar = x
rice = y
5x + 7y = 153......Eq(1)
7x + 5y = 147......Eq(2)
Multiply Eq(1) by 5
Multiply Eq(2) by 7
25x + 35y = 765....Eq(1)
49x + 35y = 1029...Eq(2)
Subtract Eq(1) from Eq(2)
24x = 264
x = 11
Substitute x = 11 into Eq(1)
5x + 7y = 153......Eq(1)
5(11) + 7y = 153
55 + 7y = 153
7y = 153 - 55
7y = 98
y = 14
x = 11; y = 14
6kg of sugar
6(11) = 66
10kg of rice
10(14) = 140
Total price = 206
Hope this helps :-)